A tow truck pulls a 1300 kg car stalled in a muddy road. What is the power of the truck if it moves the car to a distance of 5 m in 15 s? The coefficient of friction between the car tires and the road is 0.55. Assume that the acceleration of the car is 1m/s2
1
Expert's answer
2019-02-19T09:53:07-0500

The power of the track

Power=(Work done)/Time

P=W/t=(F×s)/t

The Newton’s second law states

ma=F-F_frict

So

〖F=ma+F〗_frict=ma+μmg

Finally

P=(m(a+μg)s)/t=(1300×(1+0.55×9.8)×5)/15=2336 W≈2.3 kW